OpenStudy (anonymous):

Gasoline is pouring into a cylindrical tank of radius 3 feet. When the depth of the gasoline is 4 feet, the depth is increasing at 0.2 ft/sec. How fast is the volume of gasoline changing at that instant? Include units

OpenStudy (anonymous):

i know the goal is to find dv/dt, but im not sure on how to do that \[v=\pi r^2*h\]

OpenStudy (cruffo):

first find the volume when r = 3 and h = 4 ( i got \(v=36\pi\)) Thus you know: r = 3 h = 4 v = 36\(\pi\) dr/dt = 0 (radius is not changing) dh/dt = 1/5 Take derivative of formula with respect to time (must use product rule), substitute known values, and solve for dv/dt:

OpenStudy (anonymous):

i'll try it out thanks!

OpenStudy (cruffo):

as a check, i got \(\large\dfrac{9\pi}{5}\) ft/sec.

OpenStudy (anonymous):

\[dv/dt=\pi *r^2*(dh/dt)\] and once i subsitute, i think i do get 9pi/5

OpenStudy (anonymous):

thanks alot!!

OpenStudy (cruffo):

np